f(x) = { 1 x (3-X) it osx=2 Co otherwise

Let X be a random variable with probability density function

a) Find the mean of X

b) Find the standard deviation of X round to four decimal places.

c) Let G = X2  Find the probability density function fG of G

Show work for each part plz
Answer #1

= f(x) = { / 0x (3-X) y OC XE2 0 میشلاق ) Mean of x = E(X) = Sx. f G ) di 2 x (3-x) da 194-* de (1) = [8-4] - *** - 2 Standar3 cq. 92 2 * [ 1-6 ] 1 = 20 [ 12 - 32 ] ( 60-7°) - Now. V(x) = E(X2) - [Ex] = -1 25 43 ) [from (i) fci)] Standard deviation o